Balance each of the following equations according to the half-reaction

method:
1) CN-​ ​(aq) + ClO​2(​ aq) —> CNO-​ ​(aq) + Cl-​ ​(aq) (in acid)

2) MnO​4-​ ​(aq) +NO​2-​ ​(aq) → MnO​2(​ s) + NO3​ -​ ​(aq) (in base)

3) In which species does nitrogen have the highest oxidation number? a) NaNO​3

b) HNO​2 c) NO​2-​ d) NH​3 e) N​2
Answer #1

cont clg (No tell H₂H+ oto CN I el ICNO I al *=-1 X+2(-2)0 x+(-2)27 x=+4 xat ! oxidation half reaction | Reductim half geactining + Ng ning & Not AHO o- oH - x+41-2)=- x=+7 +2(-2)=-1| X + 2(-2) 20 x+3 (-2) = x=+3 x=+4 Txas oxidation half reaction |HNO, NANO ++&+31-2)=0 Itx-620 x-520 x=+5 1+ X+2(-2)=0 1+X-420 2-320 x=+3 NH3 X+3(+1) 20 x=-3 NO- X+2(-2) = -1 x-4201 x=.4-1=+